Hello-
I have a question that I solved but I am not sure if I got the correct answer. The problem is: If two painters can finish painting a house in 7 hours, and it takes the slower painter 19 hours by himself, how long would it take the faster painter to finish the house.
Here is what I did: Let f=faster painter hours and s=slower painter hours. I have 1/f + 1/s = 1/7. But 1/s = 1/19. So 1/f = 1/7 - 1/19 = 12/133. So f= 133/12 which is about 11 hours. Is that correct? If not, please tell me where I went wrong.
Thank you very much!

Answer by ewatrrr(24785)

 
Hi,
1/f = 1/7 - 1/19 = 12/133 Yes, good Work
And 1/f = 133/12 = 11 112 hr 0r 11hr and 5min